# Investigating the Representation of Backchannels and Fillers in Fine-tuned Language Models

### v1

@@ Appendix A Data Statistics
In this section, we include details on the different backchannels/fillers used in this paper, as shown in Tables 1 and 2. The Japanese data is selected on the basis of the guidance and analysis shown in Kawamori et al. (1996) and Onishi et al. (2023). The English data is based on data descriptions in the Switchboard Corpus Holliman et al. (1992). In the main text, we selected the 15 most frequent backchannels/fillers for our data analysis. In the MapTask (Anderson et al., 1991) and Switchboard datasets, we select the following backchannels/fillers as our study objects. In the BTSJ 1000 Person Japanese Natural Conversation Corpus (BTSJ) (Usami, 2023) dataset, we use the following backchannels/fillers as our study objects. It is important to note that the backchannels/fillers without comma are used to represent its different variants, for example ‘うん’(un, ‘yeah’) includes ‘うん、’(un,) and ‘うん。’(un.).
Table 1: The top 15 selected English backchannels/fillers and their frequency in Switchboard and MapTask (combined), which contain 127672 backchannels/fillers.
@@
Table 2: The top 15 selected Japanese backchannels/fillers and their English transcription and frequency in the BTSJ corpus, which contains 170 898170\,898 backchannels/fillers.
Backchannels/Fillers Transcription Occurrence
うん (‘うんうん’, ‘うんうんうん’, ‘ううん’, ‘うーん’, ‘うんー’) ‘un’ 23.18%
はい (‘はいー’, ‘はーい’) ‘hai’ 17.41%
あ (‘ああ’, ‘あああ’, ‘あー’, ‘あっ’) ‘a’ 8.67%
え (‘ええ’, ‘えええ’, ‘えー’, ‘えっ’) ‘e’ 6.08%
そう (‘そうそう’, ‘そうそうそう’, ‘そうー’, ‘そーう’) ‘sou’ 5.25%
ま (‘まー’, ‘まあ’, ‘まあー’) ‘ma’ 4.91%
なんか (‘なんかー’, ‘なんかね’, ‘なんかねー’) ‘nanka’ 4.73%
あの (‘あのー’, ‘あのね’) ‘ano’ 4.25%
ん (‘んー’) ‘n’ 2.67%
そうです (‘そうですね’, ‘そうですねー’, ‘そうですよ’, ‘そうですよね’,
‘そうですよねー’, ‘そーうですね’, ‘そうっすね’)
‘soudesu’ 2.40%
は (‘はは’, ‘ははは’, ‘はー’, ‘はあ’, ‘はあー’, ‘はっ’) ‘ha’ 2.24%
ね (‘ねー’) ‘ne’ 2.19%
いや (‘いやいや’, ‘いやいやいや’, ‘いやー’) ‘iya’ 1.77%
@@ Appendix B Data Pre-processing
As we observed in our data, fillers are usually surrounded by others words with substantial meaning, backchannels on the other hand, usually stand out as an independent utterance. Therefore, our data preprocessing aims to contextualize the backchannels/fillers that are not surrounded by other words so that their meaning can be reflected by the other words, i.e., distributional semantics (Firth, 1968, pp. 1-32, Mikolov et al., 2013). The illustrative example which reflects our concept is Figure 1.
We noticed that some of the backchannels/fillers can be subsequences of other words (e.g., ‘um’ is a subsequence of ‘maximum’), which affects the tokenization of the data when we add the backchannels/fillers in the vocabulary as special tokens. Another case is that some of the backchannels/fillers can be ambiguous, for example, ‘well’ and ‘right’ in English can have substantial meaning as well as serve as feedback with either positive or negative emotion to the previous utterance. The same case holds for the Japanese word ‘ちょっと’ (chotto), which can either mean ‘a little’ or as a backchannels/fillers, indicate, e.g., hesitation. Our solution is to add a special token to the backchannels/fillers so that the tokenizer will not mistake them. This practice is based on our observation in the data that for the backchannels/fillers which are ambiguous, the use of these words as backchannels/fillers usually occur as the first word in the utterance, followed by a comma as a filler (e.g., ‘well,’ or co-occur with other backchannels/fillers (e.g., Okay, right).
@@ Appendix C Implementation Details
We used a total of eight L40 48G GPUs for our experiments, with one GPU assigned to each task. For the BERT model, the runtime for each task is approximately 3 hours for both English and Japanese, which includes fine-tuning and extracting embeddings for the clustering tasks. For fine-tuning GPT-2 Japanese and English, each task of the Japanese experiment takes about 2 hours, which includes fine-tuning and extracting embeddings for the clustering tasks. For English, the tasks take about 1 hours in total. As for the fine-tuning of TurnGPT tasks for Japanese and English, we used the following parameters: batch size $3$; weight decay $0.01$; dropout rate $0.1$; learning rate $0.0001$. A total of $15$ epochs were used to complete the fine-tuning tasks. After each epoch, a checkpoint (model) was generated, saving the weight parameters gained during training. The model with the minimum loss value was chosen as the final model for estimating the probability of the turn transition potential.
For LLaMA-3 8B and Qwen-3 8B, we used LoRa to accelerate the fine-tuning process. The parameters for LoRA were set as follows. The rank of the low-rank matrices is 16, and a dropout rate of 0.1 is used to the LoRA layers to improve regularization. LoRA is applied to the ‘q_proj’ and ‘v_proj’ layers within the model’s attention mechanism. For LLaMA-3 8B, each task takes about 8 hours in the Japanese data and about 7 hours in the English data. In contrast, the Qwen-3 8B model requires around 15 hours and 10 hours for the same tasks, respectively.
For $k$-means clustering, we first standardize the obtained embeddings and apply PCA (Principal Component Analysis) to reduce the dimensionality to 100, facilitating subsequent clustering operations. The number of clusters $k$ ranges from 2 to 15, and the optimal $k$ is selected based on the highest silhouette score achieved.
@@ C.1 Other Method for Obtaining Meaning Representation
Parallel to fine-tuning, there are also some other methods to acquire the meaning representations. In terms of learning the representation of backchannels/fillers in LMs, one important method is contrastive learning Gao et al. (2021). To the best of our knowledge, the only highly relevant work for our study is by Qian and Skantze (2024), who use contrastive learning methods to test how the speech models HuBert Hsu et al. (2021) and Whisper Radford et al. (2023), as well as the language model BERT can represent the different functions of feedback signals (namely backchannels). Their results show that the learnt embeddings can carry information about different functions backchannel possesses (see Figure 3 in Qian and Skantze, 2024 for details). In this paper, we focus on fine-tuning only as our method to acquire the representation of backchannels/fillers. We leave contrastive learning aside, since, unlike in Qian and Skantze (2024), the types of backchannels/fillers are much larger and thus induce much higher computational costs. The detailed reason for not using contrastive learning (performed in Qian and Skantze, 2024, while the setting and study objects are different and simpler) to get the representation of backchannels/fillers is mainly due to its computational cost and uncertainty. As an example, in the Japanese data, although we listed the most frequent backchannels/fillers in our paper, we also have the tail examples (those examples with very few occurrence, less than 50 times in the whole dataset) and together we have more than 80 types of backchannels/fillers.
The most important part of contrastive learning is its negative sampling mechanism. Negative sampling works in the way that given a positive example (in our case natural utterance with a backchannel/filler), we generate several negative samples (so some synthesised utterances where the original backchannels/fillers are replaced with random ones) and use them jointly in training so that positive examples are closer to each other. The reason this was doable and worked well in Qian and Skantze (2024) is that the candidate negative examples are just selected from the backchannels which are classified as feedback, thus limited negative sample candidates and therefore much lower computational cost. Moreover, both positive and negative samples are feedback but with different function types so the experimental results are quite controllable.
In this paper, we examine backchannels/fillers in general. If we perform contrastive learning, the situation becomes much more complex as we have to cover a lot of negative examples (a mixture of backchannels/fillers), which incurs a higher computational cost. Furthermore, we are unsure whether meaningful representations can be obtained if negative sampling is allowed to work randomly.
@@ Appendix D Further Supporting Results
D.1 Dialogue Act Classification
Our biggest concern toward the experiment results reported in this study is that, whether our fine-tuning tasks which are used to leverage the usage of backchannels and fillers in LM, will potentially undermine LM’s general performance. Therefore, we investigate this issue by checking LM’s capability of solving the dialogue act prediction task, which is feasible in Switchboard and MapTask datasets. We specifically look at, before and after fine-tuning, whether the accuracy on dialogue act prediction will change drastically or not. The results are summarised in the Table 3. As shown in Table 3, BERT, GPT-2 and Qwen3 exhibit a slight decrease in accuracy after fine-tuning, while LLaMA-3 not only maintain their performance but even show improvements. These results suggest that our fine-tuning strategy does not compromise the language understanding ability of models and may, in some cases, yield modest gains.
Model no fine-tuned fine-tuned
@@ D.1 Dialogue Act Classification
Table 3: Accuracy on the dialogue act classification task (MapTask dataset) with linear probing. We use 5 training epochs for BERT and GPT-2, and 3 training epochs for LLaMA-3 and Qwen-3.
D.2 Analysis of the Generation of Backchannels and Fillers
@@
As further evidence of improved representation capability of backchannels/fillers in fine-tuned LMs, in this subsections, we performs a qualitative analysis on the generation results under the fine_tuning_no_context and fine_tuning_one_context settings respectively. We selected around 4000 utterances from both English and Japanese corpora. As summarised in Tables 4 and 5, the results show that the fine-tuned LLaMA-3 model increases the usage of backchannels/fillers in the generation task (i.e., the frequency increases). Moreover, different types of backchannels/fillers are used (increasing diversity). This is further and crucial evidence showing that the models do learn the representation of backchannels/fillers.
Here we include four examples (Figures in  6(a)) generated by LLaMA-3-8b model in English as our qualitative analysis on the LLM’s capability of generating backchannels and filler after fine-tuning. In the examples, the tag <ds><ds> is used to indicate backchannels and fillers, /A and /B mark the speakers and their utterance. One thing we notice is that, it is under the fine_tuning_one_context setting that LLaMA-3 begins to use backchannels/fillers in its generation, which indicates the effectiveness of fine-tuning and context information. The example 1 and example 4 show that, fine-tuned LLaMA-3 can generate feedback signal (e.g., yeah) to acknowledge the previous utterance. It can also use filler to mimic disfluency or hesitation in its utterance (e.g., um in both examples). In example 2, we can see that, fine-tuned LLaMA-3 can generate a backchannel to show attentiveness to the previous speaker’s utterance (the uh-huh to the utterance by speaker A). In example 3, we can see a more complex usage of fillers: the first filler so indicates the transition of topic, and again a filler um indicates disfluency or cognitive load in the production of the speech. The examples shown here all support the idea that, we can indeed fine-tune a LM to become a conversational agent which mimic the way human beings produce speech.

### v2

@@ Acknowledgements
The early structure of this paper benefited significantly from discussions with Lívia Qian, Erik Ekstedt and Siyang Wang from TMH Department, KTH as well as group members from CLCS lab, SUSTech. We sincerely thank all the reviewers for their feedback on the paper. This work is funded by the Deutsche Forschungsgemeinschaft (DFG): TRR 318/3 2026 – 438445824, A02. This work also acknowledges the funding from Shenzhen Science and Technology Program (No. JCYJ20240813094612017) and Guangdong Province ZJRC Program (No. 2024QN11X145). This work also acknowledges the funding from the Swedish Research Council (VR) project 2020-03812.
@@ Appendix A Data Statistics
In this section, we provide details on the backchannels/fillers used in this paper, see Table 1 (English) and Table 2 (Japanese). We selected the 15 most frequent backchannels/fillers for our data analysis.
Table 1 shows the backchannels/fillers in the Switchboard (Godfrey and Holliman, 1993) and MapTask (Anderson et al., 1991) corpora that we selected as our study objects.
Table 1: The top 15 selected English backchannels/fillers and their combined frequency of occurrence in Switchboard and MapTask (from a total number of 127 672127\,672 backchannels/fillers).
@@
Table 2 shows the backchannels/fillers in the BTSJ 1000 Person Japanese Natural Conversation Corpus (BTSJ) (Usami, 2023) that we selected as our study objects. The Japanese data was selected based on prior guidance and analyses (Kawamori et al., 1996; Onishi et al., 2023). In the table, the backchannel/filler is used to represent its different variants, for example ‘うん’ (un, ‘yeah’) includes ‘うん、’ (un,) and ‘うん。’ (un), etc.
Table 2: The top 15 selected Japanese backchannels/fillers (with variants), their transcription, and frequency of occurrence in the BTSJ corpus (from a total number of 170 898170\,898 backchannels/fillers).
Backchannel/Filler Variants Transcription Occurrence
うん ‘うんうん’, ‘うんうんうん’, ‘ううん’, ‘うーん’, ‘うんー’ ‘un’ 23.18%
はい ‘はいー’, ‘はーい’ ‘hai’ 17.41%
‘ああ’, ‘あああ’, ‘あー’, ‘あっ’ ‘a’ 8.67%
‘ええ’, ‘えええ’, ‘えー’, ‘えっ’ ‘e’ 6.08%
そう ‘そうそう’, ‘そうそうそう’, ‘そうー’, ‘そーう’ ‘sou’ 5.25%
‘まー’, ‘まあ’, ‘まあー’ ‘ma’ 4.91%
なんか ‘なんかー’, ‘なんかね’, ‘なんかねー’ ‘nanka’ 4.73%
あの ‘あのー’, ‘あのね’ ‘ano’ 4.25%
‘んー’ ‘n’ 2.67%
そうです ‘そうですね’, ‘そうですねー’, ‘そうですよ’, ‘そうですよね’, ‘そうですよねー’, ‘そーうですね’, ‘そうっすね’ ‘soudesu’ 2.40%
‘はは’, ‘ははは’, ‘はー’, ‘はあ’, ‘はあー’, ‘はっ’ ‘ha’ 2.24%
‘ねー’ ‘ne’ 2.19%
いや ‘いやいや’, ‘いやいやいや’, ‘いやー’ ‘iya’ 1.77%
@@ Appendix B Data Pre-Processing
As we observed in our data, fillers are usually surrounded by others words with substantial meaning, backchannels on the other hand, usually stand out as an independent utterance. Therefore, our data preprocessing aims to contextualize the backchannels/fillers that are not surrounded by other words so that their meaning can be reflected by the other words, i.e., distributional semantics (Firth, 1957, pp. 1–32, Mikolov et al., 2013). Figure 1 illustrates our concept.
We noticed that some of the backchannels/fillers can be subsequences of other words (e.g., um is a subsequence in ‘maximum’), which affects the tokenization of the data when we add backchannels/fillers to the vocabulary as special tokens. In addition, some backchannels/fillers can be ambiguous, for example, ‘well’ and ‘right’ in English can have substantial meaning, but also serve as feedback, expressing either a positive or negative attitude towards the previous utterance. The same applies to the Japanese word ‘ちょっと’ (‘chotto’), which can mean ‘a little’ or be used as a backchannel/filler to indicate hesitation, for example. Our solution to this problem is to add a special token to backchannels/fillers so that the tokenizer will not mistake them. This practice is based on our observation in the data that for the backchannels/fillers which are ambiguous, the use of these words as backchannels/fillers usually occur as the first word in the utterance, followed by a comma as a filler (e.g., ‘well,’ or co-occur with other backchannels/fillers (e.g., okay, right).
@@ Appendix C Implementation Details
We used a total of eight L40 48G GPUs for our experiments, with one GPU assigned to each task. For the BERT model, the runtime for each task was approximately 3 hours for both English and Japanese, which includes fine-tuning and extracting embeddings for the clustering tasks. For fine-tuning GPT-2 Japanese and English, each task of the Japanese experiment took about 2 hours, which includes fine-tuning and extracting embeddings for the clustering tasks. For English, the tasks took about 1 hour.
For the fine-tuning of TurnGPT tasks for Japanese and English, we used the following parameters: batch size $4$; weight decay $0.01$; dropout rate $0.3$; learning rate $0.0005$. A total of $15$ epochs were used to complete the fine-tuning tasks. After each epoch, a checkpoint (model) was generated, saving the weight parameters gained during training. The model with the minimum loss value was chosen as the final model for estimating the probability of the turn transition potential.
For LLaMA-3 8B and Qwen-3 8B, we used LoRa to accelerate fine-tuning. The parameters for LoRA were set as follows. The rank of the low-rank matrices is $16$, and a dropout rate of $0.1$ is used to the LoRA layers to improve regularization. LoRA is applied to the ‘q_proj’ and ‘v_proj’ layers within the model’s attention mechanism. For LLaMA-3 8B, each task took about 8 hours for the Japanese data and about 7 hours for the English data. In contrast, the Qwen-3 8B model required around 15 and 10 hours for the same tasks, respectively.
For $k$-means clustering, we first standardized the obtained embeddings and applied PCA (Principal Component Analysis) to reduce the dimensionality to 100, facilitating subsequent clustering operations. The number of clusters $k$ ranged from 2 to 15, and the optimal $k$ was selected based on the highest silhouette score achieved.
@@ C.1 Other Method for Obtaining Meaning Representation
Parallel to fine-tuning, there are also some other methods to acquire the meaning representations. In terms of learning the representation of backchannels/fillers in LMs, one important method is contrastive learning Gao et al. (2021). To the best of our knowledge, the only highly relevant work for our study is by Qian and Skantze (2024), who use contrastive learning methods to test how the speech models HuBert Hsu et al. (2021) and Whisper Radford et al. (2023), as well as the language model BERT can represent the different functions of feedback signals (namely backchannels). Their results show that the learnt embeddings can carry information about different functions a backchannel possesses (see Figure 3 in Qian and Skantze, 2024 for details). In this paper, we focus on fine-tuning only as our method to acquire the representation of backchannels/fillers. We leave contrastive learning aside, since, unlike in Qian and Skantze (2024), the types of backchannels/fillers are much larger and thus induce much higher computational costs. The detailed reason for not using contrastive learning to get the representation of backchannels/fillers is mainly due to its computational cost and uncertainty (the setting and objects of study are different and simpler in Qian and Skantze, 2024). As an example, in the Japanese data, although we listed the most frequent backchannels/fillers in our paper, we also have the tail examples (those examples with very few occurrence, less than 50 times in the whole dataset) and together we have more than 80 types of backchannels/fillers.
The most important part of contrastive learning is its negative sampling mechanism. In negative sampling, given a positive example (a natural utterance with a backchannel/filler), several negative samples are generated (synthesised utterances in which the original backchannels/fillers are replaced with random ones). These negative samples are then used jointly in training to move positive examples closer to each other. In Qian and Skantze (2024) this was doable and worked well because the candidate negative examples are just selected from the backchannels which are classified as feedback, thus limited negative sample candidates and therefore much lower computational cost. Moreover, both positive and negative samples are feedback but with different function types so the experimental results are quite controllable.
@@ Appendix D Further Supporting Results
D.1 Representation of Backchannels/Fillers From LMs’ Different Hidden Layers
In the main text, we evaluate the improvement of representations by applying clustering analysis on the weights extracted from the last hidden layers of the LMs. Here we include additional experiments to further investigate the improvement from different layers before and after fine-tuning. To reflect the difference from different hidden layers, for LLaMA-3 8B, we selected layers 8, 16, 24, 32, for Qwen-3 8B, we selected layers 9, 18, 27, 36, for GPT-2 English and BERT models, we selected layers 4, 8, 12, for GPT-2 Japanese model, we selected layers 6, 12, 18, 24. The results are shown in Figure 7 and Figure 8.
Figure 7: Improvement of the representation of backchannels/fillers on selected layers. Setting: no-context.
Figure 8: Improvement of the representation of backchannels/fillers on selected layers. Setting: one-context.
First of all, it turns out that improvement of representation in different selected layers, as quantified by silhouette score, is in general observable in the fine-tuned models in both settings. Second, before fine-tuning silhouette scores are generally higher in the shallow layer than the deeper layers, which indicates that shallow layers have better representations of backchannels/fillers compared to deeper layers. Fine-tuning seems to break this tendency. For example, fine-tuned Llama-3 8b models, under two settings, have improved representation in their deeper hidden layers. Similar patterns can also been seen in Japanese BERT, GPT models and Qwen-3b for English.
D.2 General Performance of Fine-tuned LMs
Our biggest concern toward the experimental results reported in this study is whether the fine-tuning tasks, used to leverage the usage of backchannels and fillers in LMs, has the potential to undermine an LM’s general performance. We therefore investigate this issue by checking the LMs’ capability of doing a dialogue act prediction task, which is, in principle, feasible on the Switchboard and MapTask datasets. A dialogue act is an annotation label on utterance(s) of a dialogue that indicates the communicative functions of the utterance. The communicative functions include ‘statement’, ‘agreement/accept’, ‘wh-question’, ‘backchannel/acknowledge’, etc. (Stolcke et al., 2000). We specifically compare the accuracy on dialogue act prediction task before and after fine-tuning and examine whether performance changes. The results are summarised in the Table 3. As can be seen, BERT, GPT-2, and Qwen-3 exhibit a slight decrease in accuracy after fine-tuning, while LLaMA-3 shows improvements. These results suggest that our fine-tuning strategy does not compromise the language understanding ability of models and may, in some cases, even yield modest gains.
Table 3: Accuracy on the dialogue act classification task (MapTask dataset) before and after fine-tuning, with linear probing. We used five training epochs for BERT and GPT-2, and three training epochs for LLaMA-3 and Qwen-3.
Model no fine-tuning fine-tuning
@@ D.2 General Performance of Fine-tuned LMs
Dialogue 1 Input: /A <ds>yeah</ds> /A /// /B so we went in and the financed the car for five and we didn’t know that we got kind of taken just a little bit so we kind of upset that we having to you know pay this fifth year on the car because it was just not a wise thing but we learned a good lesson so you know /B /// Ground truth: /A <ds>well</ds> five years is not that <ds>uh</ds> five years is not that bad of a problem only because you paying more interest but your payments are lower /A Output (no_ft, one_ctx): /A this all the experience that we got in that situation yeah I just want to say that it was not just the car that we had to deal with but it was /A Output (ft, one_ctx): /A <ds>yeah</ds> we were just <ds>um</ds> my my parents bought the car that we would drive in high school and they put like two hundred down and and I … /A
Dialogue 2 Input: /A but then <ds>uh</ds> we also didn’t do spend <ds>uh</ds> time with them trying to trying to <ds>uh</ds> you know just family time /A /// /B <ds>uh-huh</ds>/B /// Ground truth: /A <ds>uh</ds> we had a at a fairly young age we started teaching them to play games and card games and so forth and and <ds>uh</ds> <ds>uh</ds> but we didn’t <ds>uh</ds> do it as much or as often as <ds>uh</ds> I would have liked to I would like to do that as a as a family and /A Output (no_ft, one_ctx): /A with the family that we have now with the family that we have now. What are you going to do with that kind of logic? I mean, I’d like to solve this problem. But I can’t. Can you help me? I have a little bit of time before the family comes over. So, can you give … /A Output (ft, one_ctx): /A I think <ds>uh</ds> my father was really into them and we would you know go camping and stuff and <ds>uh</ds> you know I think if you do that then it important to spend some time with them you know /A /// /B <ds>uh-huh</ds> /B
Dialogue 3 Input: /A so it becomes <ds>uh</ds> a general battle of the w- sort of the election of the lesser of two evils I think so /A /// /B <ds>yeah</ds> /B /// Ground truth: /A I don’t know I just <ds>uh</ds> I unhappy with it but and well I I sort of semi-fortunate right now I a graduate student so I don’t make that much /A Output (no_ft, one_ctx): /A The concept of "lesser of two evils" suggests that when faced with a difficult decision, it’s better to choose the option that is less undesirable. In the context of the 202 /A Output (ft, one_ctx): /A it always seems to be that way <ds>so</ds> <ds>um</ds> what do you think is the other major issue now /A /// /B <ds>um</ds> I I don’t know I think …/B
Dialogue 4 Input: /A <ds>yeah</ds> <ds>well</ds> my <ds>uh</ds> my <ds>uh</ds> probably one of the biggest decisions I think that was very strengthening for our family was rather than have one child make that decision /A /// /B <ds>uh-huh</ds> /B /// Ground truth: /A than just delegate it I think that they they had a great deal of <ds>um</ds> all the brothers and sisters got together and they actually had a conference /A Output (no_ft, one_ctx): /A one year. For one of one or a rather long time I am one I do one I did one I have one I will one one I had one /A Output (ft, one_ctx): /A I think we went to church and I think we sought <ds>um</ds> the advice of those that could see it from a different perspective /A
Dialogue 5 Input: /A you don’t have a stone circle? /A /// /B no /B /// /A <ds>okay</ds> do you have a stone circle on the map at all /A /// /B <ds>um</ds> i have a ruined monastery /B /// Ground truth: /A fine <ds>um</ds> do you have any obstacle below the telephone kiosk /A Output (no_ft, one_ctx): /A hat’s good too we can explore it 1a /A Output (ft, one_ctx): /A you have a ruined monastic <ds>uh-huh</ds> <ds>well</ds> you’re going to go down and then you’re going to go right and then you’re going to go up over the top of the stone monument and then down again /A
Dialogue 6 Input: /A <ds>eh</ds> can you dr– actually draw round the /A /// /B <ds>well</ds> <ds>yeah</ds> i suppose /B /// /A <ds>well</ds> if you could draw round them then/A /// Ground truth: /B straight down to the giraffes /B Output (no_ft, one_ctx): /B then maybe we can make a plan 1. but then i can’t actually get from the other end of the rope bridge to the giraffe /B Output (ft, one_ctx): /B i don’t know whether there’s a river or not i’ve got a river which is on the left-hand side /B
Figure 9: Six illustrative examples generated with the non fine-tuned and fine-tuned LLaMA-3 model in the one-context setting.
D.3 Further Qualitative and Quantitative Analyses of Generation
@@
As further evidence of improved representation capability of backchannels/fillers in fine-tuned LMs, in this subsections, we performs a qualitative analysis on the generation results under the fine_tuning, no_context and fine_tuning, one_context settings respectively. Both fine-tuned and non fine-tuned LMs are guided to perform an NLG task which requires them to complete the dialogue based on the given context. We selected around 4000 utterances from both English and Japanese corpora. As summarised in Tables 4 and 5, the results show that the fine-tuned LLaMA-3 model increases the usage of backchannels/fillers in the generation task (i.e., the frequency increases). Moreover, different types of backchannels/fillers are used (diversity increases). This is further and crucial evidence showing that the models do learn the representation of backchannels/fillers.
As our small scale human evaluation, here we select six representative dialogue examples (see Figure 9) generated by the LLaMA-3 8b model in English as our qualitative analysis on the LLM’s capability of generating backchannels and fillers after fine-tuning. Below is the instruction for reading the dialogue examples:
  • /A…/A and /B…/B: the beginning and the end of the utterance(s) from speaker A, speaker B.
  • ///: marking of turn shifts.
  • <ds>…</ds>: annotation of backchannels/fillers.
  • Input: the incomplete dialogues used for the generation task.
  • Ground truth: the utterances which are the continuation of the incomplete dialogue. (Input)
  • Output (no_ft, one_ctx): The generated utterances from the off-the-shelf LLaMa-3 model under the one-context setting.
  • Output (ft, one_ctx): The generated utterances from the fine-tuned LLaMa-3 model under the one-context setting.
We notice that only under the ft, one_cx setting (fine-tuned, one-context ) does the LLaMA-3 model begin to use backchannels/fillers in its generated output, which further indicates the effectiveness of fine-tuning and context information.
The ft, one_ctx outputs of examples dialogues 1 and 4 show that the fine-tuned LLaMA-3 model can generate feedback signals (e.g., yeah) to acknowledge the previous utterance. It can also use fillers to mimic disfluency or hesitation in an utterance (e.g., um in both examples).
In output (ft, one_ctx) of example dialogue 2, we can see that the fine-tuned LLaMA-3 model can generate a backchannel to show attentiveness to the previous speaker’s utterance (the uh-huh in response to the utterance by speaker A).
In output (ft, one_ctx) of dialogue example 3, we can see a more complex use of fillers: the first filler so indicates the transition of topic, and again a filler um indicates disfluency or cognitive load in the production of speech. The examples shown here all support the idea that we can indeed fine-tune a LM to become a conversational agent which mimics the way humans produce speech.
A notable finding is the emergence of advanced pragmatic competencies of backchannels and fillers in the fine-tuned LMs. As illustrated in output (ft, one_ctx) of dialogue examples 3 and 5, the fine-tuned LLaMA-3 model successfully distinguishes between the structural role of so, a filler which marks a sequence transition, and the cognitive signalling of the filler um.
Similarly, in output (ft, one_ctx) in example dialogue 5, the uh-huh is a backchannel which indicates confirmation and signals grounding while well is a filler which shows transition and framing. Furthermore, as evidenced by example dialogue 6, the fine-tuned LM does not overgenerate backchannels and fillers to complete every dialogue. Instead, it demonstrates a nuanced capacity to omit backchannels and fillers when they are not pragmatically required.
D.4 Supporting Details for NLG Evaluation Result Shown in Figure 6